CALCULATE THE DISTANCE BETWEEN SUCCESSIVE POSITION OF THE MOVABLE MIRROR OF MICHELSONS INTERFEROMETER GIVING BEST FRINGES OF SODIUM SOURCE HAVING LINES 5896 A° AND 5890 A°.


1
Expert's answer
2021-05-24T14:13:06-0400

"\lambda_1=5890, \lambda_2=5896"


As Difference in wavelength-


"\Delta \lambda=\lambda_1-\lambda_2=\dfrac{\lambda_1 \lambda_2}{2(x_2-x_1)}"


"(x_2-x_1)=\dfrac{\lambda_1 \lambda_2}{2\Delta \lambda}"


"=\dfrac{ 5.896\times 10^{-7} \times 5.890\times 10^{-7}}{2(5896-5890}"


"(x_2-x_1)=28.94 A"